# TabulaX: Leveraging Large Language Models for Multi-Class Table Transformations

### v1

@@ V-C Performance of classification
V-D Performance Compared to baseline models
TABLE III: Performance compared to the baselines
TabulaX DTT GXJoin AFJ Ditto
@@ V-D Performance Compared to baseline models
To benchmark our approach, we compare TabulaX against four state-of-the-art baselines: Deep Tabular Transformer (DTT) [2], GXJoin [3], Auto-FuzzyJoin (AFJ) [6], and Ditto [29]. DTT leverages language models to generate outputs directly from provided examples, focusing on string transformations. GXJoin extends the CST method [9] and exhaustively searches for general textual string-based transformations to facilitate table joinability. AFJ employs similarity functions to identify the most probable rows to join without generating explicit transformation functions. Ditto is an entity matching framework that fine-tunes pre-trained language models (specifically, DistilBERT [52] in our experiments) for matching entities across different tables.
@@
Table III summarizes the performance of our framework, TabulaX, and the baselines. As shown, TabulaX consistently achieves higher or comparable F1-scores across all datasets compared to these baselines. On datasets dominated by string-based transformations, such as WT and SS, the performance gap between TabulaX and DTT is relatively small, with DTT being the best-performing baseline in our setup. Specifically, TabulaX achieves an F1-score of 0.983 on the WT dataset, only marginally higher than DTT’s 0.950. On the SS dataset, both methods perform almost identically, with F1-scores of 0.952. This similarity is expected since both models handle string transformations effectively, though TabulaX offers the added advantage of generating interpretable transformation functions. In contrast, DTT directly outputs target values without transparency, which can be a limitation in sensitive environments and scenarios requiring interpretability. The performance gap of TabulaX and the baselines increases on the KBWT dataset, which includes more complex cases such as numeric and general knowledge transformations. Here, TabulaX achieves an F1-score of 0.567, significantly outperforming DTT, which only scores 0.254. The gap highlights TabulaX’s versatility and its ability to generalize across diverse data types, a feature not shared by the string-based baselines. Additionally, TabulaX’s interpretable transformations are valuable in understanding and verifying the mappings, unlike DTT’s black-box approach.
@@
GXJoin, performs an exhaustive search in the string-based transformation space to generate interpretable transformations and is the best-performing state-of-the-art approach generating an interpretable mapping in our experiments. This large search space allows GXJoin to achieve high precision on all datasets, as it applies transformations without relying on edit distance-based matching. However, the recall is noticeably lower, indicating that it fails to transform a significant portion of the data. In contrast, TabulaX maintains a high recall by utilizing a broader set of transformation units and providing a clear mapping between input rows and transformations. Similar to other baselines, GXJoin is only limited to string-based transformations and is significantly outperformed by TabulaX and DTT on KBWT dataset.
Even without employing edit distance matching, as indicated in Table IV, TabulaX outperforms GXJoin due to its ability to handle a wider variety of transformations and its use of LLMs to interpret and apply complex patterns. GXJoin’s exhaustive search methodology, while thorough, can lead to scalability issues. Since it considers all possible combinations of transformation units, its runtime can be significantly higher compared to TabulaX, especially when dealing with longer rows, larger tables, or when many transformation units are added. This makes TabulaX a more efficient and scalable solution for real-world applications. Moreover, TabulaX provides a clear mapping between input rows and the applied transformations. GXJoin’s exhaustive search does not offer such mapping, and it may apply multiple transformations without a straightforward way to trace back the exact steps for each row.
@@
In terms of runtime, for a table with $n$ rows, TabulaX requires only a single LLM call to generate transformations for string-based tables, two LLM calls (one for relationship tagging and another for transformation generation) for algorithmic tables, and $n+1$ LLM calls for general tables. In contrast, DTT, while effective primarily for string-based tables, requires $k\times n$ LLM calls, where $k$ represents the number of self-consistency trials (set to 5 in the experiments). Both TabulaX (in the worst case) and DTT have a linear growth in the LLM calls with respect to input size, and their LLM calls can be executed in parallel if resources allow. On the other hand, GXJoin has quadratic growth in runtime due to its exhaustive search methodology. This approach also has more constraints on parallel execution, making GXJoin less scalable compared to TabulaX for larger datasets.
AFJ and Ditto rely on similarity functions to find matching rows and do not generate explicit transformation functions. This limitation reduces their applicability in scenarios where understanding the transformation is essential. Additionally, they may end up with many false positives and their performance on all datasets is inferior to that of TabulaX, with particularly poor results on the KBWT dataset.
@@
TABLE IV: The framework performance under various classification and matching methods
@@
LLM Web Tables 0.985 0.980 0.983 1.501 0.068 1.000 0.729 0.816 5.497 0.269
SpreadSheet 0.955 0.949 0.952 1.944 0.094 1.000 0.818 0.828 4.879 0.180
@@
Golden Web Tables 0.975 0.971 0.973 2.223 0.094 1.000 0.690 0.770 6.362 0.308
SpreadSheet 0.953 0.949 0.951 2.132 0.104 1.000 0.798 0.809 5.270 0.202
@@
TABLE V: Performance Metrics per Transformation Class
@@
String 139 0.958 0.954 0.956 2.152 0.102 1.000 0.774 0.800 5.513 0.225
Numbers 3 1.000 1.000 1.000 - - 1.000 1.000 1.000 - -
Algorithmic 3 1.000 0.667 0.667 3.333 0.333 1.000 0.667 0.667 3.333 0.333
General 73 0.699 0.521 0.559 7.029 0.371 0.973 0.413 0.498 7.608 0.511
Table IV presents a detailed comparison of the framework’s performance under two different matching scenarios: edit-distance-based matching (left panel) and exact matching (right panel). The former matching strategy is more suited for applications where a set of target values is available, such as in unequal table joins, where small variations between the generated and expected values can be tolerated. The latter, exact matching, is more appropriate for tasks like missing data imputation, where precise outputs are required. The top panel of the table reflects results when the input table classification is determined by the LLM-based model (GPT-4o in this set of experiments), while the bottom panel provides performance metrics when using the ground truth classification, referred to as the “Golden” classifier. This comparison allows us to assess the impact of the classification on the overall performance of the framework.
An interesting observation is that the performance of the model on datasets consisting of string-based transformations (Web Tables and Spreadsheet) is slightly better when the classification is done via the LLM-based model compared to the golden classifier. For instance, on the Web Tables dataset using edit-distance matching, the F1-score with LLM-based classification is 0.983, marginally higher than the 0.973 achieved with the golden classifier. This counterintuitive performance can be attributed to the LLM-based classifier occasionally mislabeling some difficult string-based transformations as general transformations. In these challenging cases, treating them as general transformations allows the model to generate the target values directly by leveraging the LLM’s generative capabilities, without being constrained to produce a specific transformation function covering all input rows. While this approach can lead to better target predictions and higher recall, it sacrifices interpretability since the general transformations are not explicitly defined. In contrast, the golden classifier strictly labels these cases as string-based transformations, requiring the model to generate explicit transformation functions, which may be more difficult for complex patterns.
This performance increase is more noticeable on string-based datasets, where minor inconsistencies are common. Specifically, on the Web Tables dataset under LLM-based classification, the F1-score improves from 0.816 with exact matching to 0.983 with edit-distance matching, representing a substantial increase of approximately 20.5%. This demonstrates that allowing for slight variations in matching can significantly enhance the model’s ability to correctly map source to target values in real-world, noisy datasets with inherent inconsistencies. On the other hand, on the KBWT dataset, which includes more general transformations and relies on external knowledge, the improvement in F1-score when using edit-distance matching is less pronounced. Specifically, the F1-score under LLM-based classification improves from 0.512 with exact matching to 0.567 with edit-distance matching, an increase of approximately 10.7%. This indicates that for general transformations, minor discrepancies are less probable to occur when the model is not constrained to a string-based transformation to cover all rows.
V-E Class-Wise Performance and Error Analysis
@@
In this section, we analyze the model’s performance across each transformation class and highlight specific challenges encountered in some cases. This evaluation helps identify patterns in model performance and limitations that may inform future improvements. Table V provides an overview of the results for each class, with support indicating the number of tables in each category.
@@
In the string transformation class, the framework demonstrates strong performance. When edit-distance matching is applied, only a few cases are missed, primarily due to noisy data or substantial variations in formatting that were not adequately represented in the provided examples. Edit-distance matching allows for minor discrepancies, improving recall, though it occasionally results in missed rows when format variations exceed the threshold for an approximate match. When exact matching is employed, the model’s recall decreases in the string class. This drop is expected, as exact matching is more stringent and does not allow for minor discrepancies or transformations that partially transform the source into target. For instance, in a dataset containing information about state governors formatted as Name; (birth date - death date) in the source and Name - (in office years) in the target, the model successfully maps names but the date transformation is not feasible, which will adversely affect the performance when exact matching is employed.
@@
TABLE VI: Performance of the framework with Smaller and Open Source LLMs
@@
GPT-4o Web Tables 0.975 0.971 0.973 2.223 0.094 1.000 0.690 0.770 6.362 0.308
SpreadSheet 0.953 0.949 0.951 2.132 0.104 1.000 0.798 0.809 5.270 0.202
@@
GPT-4o.m Web Tables 0.924 0.920 0.922 3.752 0.142 0.983 0.626 0.685 8.214 0.372
SpreadSheet 0.979 0.976 0.977 1.763 0.072 1.000 0.802 0.806 4.730 0.198
@@
LLaMA Web Tables 0.822 0.820 0.821 6.837 0.275 0.994 0.437 0.491 11.298 0.568
SpreadSheet 0.915 0.912 0.913 3.551 0.235 0.981 0.389 0.398 8.942 0.611
@@
For numeric transformations, the model performs exceptionally well across all tables, achieving perfect precision and recall. This success can be attributed to the straightforward nature of numeric relationships, where mathematical functions can precisely capture the transformation patterns. In the algorithmic transformation class, the model maintains perfect precision but exhibits a lower recall of 0.667 in both matching strategies. The primary challenge here is the complexity of certain algorithms. For example, transforming Gregorian dates to Hijri dates involves a sophisticated algorithm that the model partially captures. While it correctly identifies the relationship among source and target, the generated code may is incomplete, leading to some missed transformations.
@@
V-F Evaluation of Different Model Sizes and Open-Source Alternatives
To evaluate how the size and type of language models affect our framework’s performance, we experimented with different commercial and open-source models of varying sizes. Table VI summarizes the results of these experiments. In this table, GPT-4o refers to the gpt-4o-2024-05-13 model, one of the leading state-of-the-art commercial models. GPT-4o.m denotes the smaller and more cost-effective gpt-4o-mini-2024-07-18 model. The LLaMA model represents the Llama-3.1-8B-Instruct, an open-source model with 8 billion parameters executable on a single GPU.
@@ V-F Evaluation of Different Model Sizes and Open-Source Alternatives
VI Conclusion and Future works

### v2

@@ 5.3. Performance of classification
5.4. Performance Compared to baseline models
Table 3. Performance compared to the baselines
TabulaX DTT GXJoin AFJ Explain-Da-V
@@ 5.4. Performance Compared to baseline models
To benchmark our approach, we compare TabulaX against four state-of-the-art baselines: Deep Tabular Transformer (DTT) (Dargahi Nobari and Rafiei, 2024), GXJoin (Omidvartehrani et al., 2024), Auto-FuzzyJoin (AFJ) (Li et al., 2021), and Explain-Da-V (Shraga and Miller, 2023). DTT leverages language models to generate outputs directly from provided examples, focusing on string transformations. GXJoin extends the CST method (Dargahi Nobari and Rafiei, 2022) and exhaustively searches for general textual string-based transformations to facilitate table joinability. AFJ employs similarity functions to identify the most probable rows to join without generating explicit transformation functions. Explain-Da-V performs example-driven structural transformations by exploring a graph of intermediate table states with an A*-based heuristic search and type-aware operators, aiming to reshape source tables so they align with the target structure.
@@
Table 3 summarizes the performance of our framework, TabulaX, and the baselines. As shown, TabulaX consistently achieves higher or comparable F1-scores across all datasets compared to these baselines. On datasets dominated by string-based transformations, such as WT and SS, the performance gap between TabulaX and DTT is relatively small, with DTT being the best-performing baseline in our setup. Specifically, TabulaX achieves an F1-score of 0.983 on the WT dataset, only marginally higher than DTT’s 0.950. On the SS dataset, both methods perform almost identically, with F1-scores of 0.952. This similarity is expected since both models handle string transformations effectively, though TabulaX offers the added advantage of generating interpretable transformation functions. In contrast, DTT directly outputs target values without transparency, which can be a limitation in sensitive environments and scenarios requiring interpretability. The performance gap between TabulaX and the baselines increases on datasets involving more complex transformations. On the TT dataset, which includes a mix of string-based, numerical, and algorithmic transformations, TabulaX achieves an F1-score of 0.918, compared to 0.643 for DTT, 0.223 for GXJoin, and 0.219 for Explain-Da-V. The TDE approach (He et al., 2018a) was proposed in the same work that introduced the TT dataset and reports a coverage of 72% on it, whereas our method achieves a recall of 92% on this benchmark. On the KBWT dataset, which contains a higher proportion of general transformations requiring external knowledge, TabulaX reaches an F1-score of 0.567, whereas DTT scores only 0.254. The gap highlights TabulaX’s versatility and its ability to generalize across diverse data types, a feature not shared by the string-based baselines. Additionally, TabulaX’s interpretable transformations are valuable in understanding and verifying the mappings, unlike DTT’s black-box approach.
@@
GXJoin, performs an exhaustive search in the string-based transformation space to generate interpretable transformations and is the best-performing state-of-the-art approach generating an interpretable mapping in our experiments. This large search space allows GXJoin to achieve high precision on all datasets, as it applies transformations without relying on edit distance-based matching. However, the recall is noticeably lower, indicating that it fails to transform a significant portion of the data. In contrast, TabulaX maintains a high recall by utilizing a broader set of transformation units and providing a clear mapping between input rows and transformations. Similar to other baselines, GXJoin is only limited to string-based transformations and is significantly outperformed by TabulaX and DTT on TT and KBWT datasets. Even without employing edit distance matching, as indicated in Table 4, TabulaX outperforms GXJoin due to its ability to handle a wider variety of transformations and its use of LLMs to interpret and apply complex patterns. GXJoin’s exhaustive search methodology, while thorough, can lead to scalability issues. Moreover, while TabulaX provides a clear mapping between input rows and the applied transformations, GXJoin’s exhaustive search does not offer such mapping, and it may apply multiple transformations without a straightforward way to trace back the exact steps for each row.
@@
In terms of runtime, for a table with $n$ rows, TabulaX requires only a single LLM call to generate transformations for string-based tables, two LLM calls (one for relationship tagging and another for transformation generation) for algorithmic tables, and $n+1$ LLM calls for general tables when lookups rely on LLM calls. Costs can be reduced if bridge tables are accessible from alternative sources. In contrast, DTT, while effective primarily for string-based tables, requires $k\times n$ LLM calls, where $k$ represents the number of self-consistency trials (set to 5 in the experiments). Both TabulaX (in the worst case) and DTT have a linear growth in the LLM calls with respect to input size, and their LLM calls can be executed in parallel if resources allow. On the other hand, GXJoin has quadratic growth in runtime due to its exhaustive search methodology. This approach also has more constraints on parallel execution, making GXJoin less scalable compared to TabulaX for larger datasets.
AFJ relies on similarity functions to identify matching rows and does not generate explicit transformation functions, which limits its applicability in scenarios where understanding the transformation process is essential. Explain-Da-V, while capable of producing interpretable transformations, is primarily optimized for structural table reshaping. Its support for value-level transformations is limited, which hinders its performance in our setup. As a result, both methods underperform compared to TabulaX across all datasets, with particularly poor results on more complex benchmarks such as TT and KBWT.
@@
Table 4. The framework performance under various classification and matching methods
@@
LLM
Web Tables (WT) 0.985 0.980 0.983 1.501 0.068 1.000 0.729 0.816 5.497 0.269
SpreadSheet (SS) 0.955 0.949 0.952 1.944 0.094 1.000 0.818 0.828 4.879 0.180
Table Transformation (TT) 0.904 0.895 0.898 2.678 0.219 0.966 0.736 0.756 4.252 0.340
@@
Golden
Web Tables (WT) 0.975 0.971 0.973 2.223 0.094 1.000 0.690 0.770 6.362 0.308
SpreadSheet (SS) 0.953 0.949 0.951 2.132 0.104 1.000 0.798 0.809 5.270 0.202
Table Transformation (TT) 0.919 0.918 0.918 2.466 0.219 0.992 0.733 0.756 3.915 0.344
@@
5.5. Impact of Matching Strategies
Table 4 presents a detailed comparison of the framework’s performance under two different matching scenarios: edit-distance-based matching (left panel) and exact matching (right panel). The former matching strategy is more suited for applications where a set of target values is available, such as in unequal table joins, where small variations between the generated and expected values can be tolerated. The latter, exact matching, is more appropriate for tasks like missing data imputation, where precise outputs are required. The top panel of the table reflects results when using the LLM-based classification model (GPT-4o in this set of experiments), while the bottom panel provides performance metrics when using the ground truth classification, referred to as the “Golden” classifier. This comparison allows us to assess the impact of the classification on the overall performance of the framework.
An interesting observation is that the model performance on datasets consisting of string-based transformations (Web Tables and Spreadsheet) is slightly better when the classification is done via the LLM-based model compared to the golden classifier. For instance, on the Web Tables dataset using edit-distance matching, the F1-score with LLM-based classification is 0.983, marginally higher than the 0.973 achieved with the golden classifier. This counterintuitive performance can be attributed to the LLM-based classifier occasionally mislabeling some difficult string-based transformations as general transformations. In these challenging cases, treating them as general transformations allows the model to generate the target values directly by leveraging the LLM’s generative capabilities, without being constrained to produce a specific transformation function covering all input rows. While this approach can lead to better target predictions and higher recall, it sacrifices interpretability since the general transformations are not explicitly defined. In contrast, the golden classifier strictly labels these cases as string-based transformations, requiring the model to generate explicit transformation functions, which may be more difficult for complex patterns.
This performance increase is more noticeable on string-based datasets, where minor inconsistencies are common. Specifically, on the Web Tables dataset under LLM-based classification, the F1-score improves from 0.816 with exact matching to 0.983 with edit-distance matching, representing a substantial increase of approximately 20.5%. This demonstrates that allowing for slight variations in matching can significantly enhance the model’s ability to correctly map source to target values in real-world, noisy datasets with inherent inconsistencies. On the other hand, on the TT and KBWT datasets, which include more algorithmic and general transformations and may rely on external knowledge, the improvement in F1-score when using edit-distance matching is less pronounced. For example, on the KBWT dataset, the F1-score under LLM-based classification improves from 0.512 with exact matching to 0.567 with edit-distance matching, an increase of approximately 10.7%. This indicates that for general transformations, minor discrepancies are less probable to occur when the model is not constrained to a string-based transformation to cover all rows.
Table 5. Performance Metrics per Transformation Class
@@ 5.5. Impact of Matching Strategies
String 207 0.957 0.953 0.955 1.795 0.120 1.000 0.775 0.802 4.144 0.223
Numbers 60 0.971 0.971 0.971 - - 0.971 0.971 0.971 - -
Algorithmic 105 0.874 0.865 0.865 3.342 0.263 1.000 0.577 0.604 5.488 0.422
General 76 0.702 0.530 0.567 6.926 0.365 0.974 0.419 0.507 7.564 0.506
5.6. Class-Wise Performance and Error Analysis
@@
In this section, we analyze the model’s performance across each transformation class and highlight specific challenges encountered in some cases. This evaluation helps identify patterns in model performance and limitations that may inform future improvements. Table 5 provides an overview of the results for each class, with support indicating the number of tables in each category.
@@
In the string transformation class, the framework demonstrates strong performance. When edit-distance matching is applied, only a few cases are missed, primarily due to noisy data or substantial variations in formatting that were not adequately represented in the provided examples. Edit-distance matching allows for minor discrepancies, improving recall, though it occasionally results in missed rows when format variations exceed the threshold for an approximate match. When exact matching is employed, the model’s recall decreases in the string class. This drop is expected, as exact matching is more stringent and does not allow for minor discrepancies or transformations that partially transform the source into target. For instance, in a dataset containing information about state governors formatted as Name; (birth date - death date) in the source and Name - (in office years) in the target, the model successfully maps names but the date transformation is not feasible via string operations, which will adversely affect the performance when exact matching is employed.
For numeric transformations, the model performs exceptionally well across all tables, achieving near-perfect precision and recall. This success can be attributed to the straightforward nature of numeric relationships, where mathematical functions can precisely capture the transformation patterns. The few cases where the model did not perform well were mostly outliers or instances where no meaningful relationship existed between the input and output values. In the algorithmic transformation class, while the overall performance remains strong, we observe a slight drop in both precision and recall. The primary challenge here is the complexity of certain algorithms. For example, transforming Gregorian dates to Hijri dates involves a sophisticated algorithm that the model partially captures. While it correctly identifies the relationship among source and target, the generated code may is incomplete, leading to some missed transformations.
@@
Table 6. Performance of the framework with Smaller and Open Source LLMs
@@
GPT-4o
Web Tables (WT) 0.975 0.971 0.973 2.223 0.094 1.000 0.690 0.770 6.362 0.308
SpreadSheet (SS) 0.953 0.949 0.951 2.132 0.104 1.000 0.798 0.809 5.270 0.202
Table Transformation (TT) 0.919 0.918 0.918 2.466 0.219 0.992 0.733 0.756 3.915 0.344
@@
GPT-4o.m
Web Tables (WT) 0.924 0.920 0.922 3.752 0.142 0.983 0.626 0.685 8.214 0.372
SpreadSheet (SS) 0.979 0.976 0.977 1.763 0.072 1.000 0.802 0.806 4.730 0.198
Table Transformation (TT) 0.906 0.904 0.905 5.325 0.273 0.990 0.671 0.695 5.002 0.423
@@
LLaMA
Web Tables (WT) 0.822 0.820 0.821 6.837 0.275 0.994 0.437 0.491 11.298 0.568
SpreadSheet (SS) 0.915 0.912 0.913 3.551 0.235 0.981 0.389 0.398 8.942 0.611
Table Transformation (TT) 0.815 0.811 0.812 7.422 0.505 0.992 0.452 0.465 7.531 0.717
@@
5.7. Evaluation of Different Model Sizes and Open-Source Alternatives
To evaluate how the size and type of language models affect our framework’s performance, we experimented with different commercial and open-source models of varying sizes. Table 6 summarizes the results of these experiments. In this table, GPT-4o refers to the gpt-4o-2024-05-13 model, one of the leading state-of-the-art commercial models. GPT-4o.m denotes the smaller and more cost-effective gpt-4o-mini-2024-07-18 model. The LLaMA model represents the Llama-3.1-8B-Instruct, an open-source model with 8 billion parameters executable on a single GPU.
@@ 5.7. Evaluation of Different Model Sizes and Open-Source Alternatives
6. Conclusion and Future works
@@
Acknowledgments
This research was partially supported by the Natural Sciences and Engineering Research Council of Canada.